OpenStudy (beccab003):

Can someone please help me with an easy question? I will medal the person with the best answer. Thanks! Find the coordinates of the ordered pair where the minimum value occurs for the equation l = 0.5x – 0.3y, given these constraints: x – y >(or equal to) -2 x + y >(or equal to) -4 2x – y <(or equal to) 4 (6, 8) (-3, -1) (0, 3) (0, -4) I believe it is (6,8). Am I right?

OpenStudy (beccab003):

@hartnn

hartnn (hartnn):

why not (-3, -1) ?

OpenStudy (amistre64):

ideally the constraints form a shape that has corners. the critical points to test will be at one of these corners.

OpenStudy (beccab003):

Because the first one is = 0.6 while (-3,-1) = -1.2. It asks for the minimum value.

hartnn (hartnn):

and 0,3 is out as it does not satisfy 1st constraint -1.2 is less than 0.6

hartnn (hartnn):

choose the least between -1.2, 0.6, 1.2

OpenStudy (beccab003):

oh...sorry I forgot about the fact that it was negative. so it is -1.2 which is (-3,-1)?

hartnn (hartnn):

yes

OpenStudy (beccab003):

Thanks! Can you check one more problem for me?

hartnn (hartnn):

sure

OpenStudy (beccab003):

Considering these two equations: y – 3x = 4 3y + 2x = 1 Which expression can be substituted for y in the bottom equation of the system to solve the system by substitution? 2/3x + 1 x = 4 – 3x 3x + 4 1 – 2x It is 3x+4

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (beccab003):

Thanks!

hartnn (hartnn):

welcome ^_^